Enter your answer in the provided box. The pH of a sodium acetate-acetic acid buffer is 4.80. Calculate the ratio [CH3COO−] / [CH3COOH].
Answer #1

Given, pH = 4.80

Now we know from the ka value tables the ka value for acetic acid = 1.8×10-5

So pKa = -log ka

   pKa = -log (1.8×10-5)

                =4.74

Now using handerson-hasselbalch equation,

           pH= pKa+ log     

           pH - pKa = log     

Putting values we get,

4.80 - 4.74   = log     

0.06 =log      

Or taking antilog , we get

      = 100.06  = 1.14

So the ratio is,

      = 1.14